# Three-dimensional quantum criticality

### v2

@@ Abstract
The SC3 phase of UTe2, which persists to $\mu_{0}H\gtrapprox$ 70 T despite its relatively modest $T_{c}\sim$ 2 K, has posed a formidable theoretical challenge to our understanding of high-$H$ superconductivity [15, 16]. As prior high-$H$ measurements in the $b-c$ plane only observed SC3 inside the PPM, the field-induced compensation of magnetic exchange [43] was proposed as a possible microscopic origin [31]. However, our finding of the zero resistance SC3 state extending outside the PPM (Fig. 4) – accompanied by an extended field region of the onset of SC3 even further away from the MM phase transition line – argues strongly against this scenario being the sole mechanism.
@@
where $i,j,k=a,b,c$ and $\chi^{-1}_{i},\beta_{ij}$, and $\delta_{ijk}$ are model parameters (see Appendix B[44, 45, 38]. The metamagnetic phase transition occurs when the free energy has two local minima at $\mathbf{M}=\textbf{M}_{0}$ and $\mathbf{M}=\mathbf{M}_{*}$ corresponding to the normal paramagnetic and PPM phases, respectively. With $\beta_{bb}<0$, the $H_{b}$ component of the external field tends to lower the minimum at $\mathbf{M}_{*}$ faster than the minimum at $\mathbf{M}_{0}$, resulting in the MM transition when $\mathcal{F}(\mathbf{M}_{*})=\mathcal{F}(\mathbf{M}_{0})$, which generically defines a 2D surface in the 3D space of $\mathbf{H}$.
@@
In general, the model parameters are temperature dependent, such that as temperature increases the two minima merge at a critical end point. For simplicity, here we neglect this temperature dependence but consider the effects of $H_{a}$ and $H_{c}$. Due to the mixed terms $\beta_{ij}$ and $\delta_{ijk}$, both $H_{a}$ and $H_{c}$ (which induce $M_{a}$ and $M_{c}$) in effect indirectly modify the $\chi_{b}$ and $\beta_{bb}$ coefficients and so act in a similar way to the temperature. As a result, large $H_{a}$ and $H_{c}$ tend to move $\mathbf{M}_{0}$ and $\mathbf{M}_{*}$ towards each other, until they merge at the critical end line (at finite temperature) or the QCL (at zero temperature) bounding the MM phase transition surface.
@@
We confirm this general picture by numerically minimizing the free energy of Eqn. (2), determining the MM transition surface and the QCL as a function of H, as shown in Figure 8. With the given parameters (see Appendix B) the model matches the experimental data from Fig. 1 reasonably well and captures the main qualitative features. Assuming SC3 is driven by the fluctuations in the vicinity of the QCL, the shape of the QCL explains most of the field direction dependence of SC3, with orbital pair breaking and spin effects potentially accounting for the rest. The QCL fluctuation scenario is further supported by the fact that, analogous to the case of ferromagnetic uranium-based superconductors such as URhGe [12, 13, 45], SC3 should be expected to occur on both sides of the QCL as we observe experimentally (Fig. 4).
Our finding of a three-dimensional quantum critical phase boundary anchoring the extremely high magnetic field superconducting phase landscape of UTe2 poses the question of what the microscopic properties of the intriguing SC3 state may be. Unlike the lower field SC2 phase – the angular domain of which has been found to exhibit a pronounced sensitivity to crystalline disorder [36] – the SC3 state has been observed even in highly disordered crystals that are too dirty to manifest SC1 [40]. Furthermore, in comparing with prior reports [31, 40] we find that the angular extent of SC3 in the $bc$ rotation plane appears to extend to very similar inclination angles irrespective of crystalline quality (see Figure 17). We speculate that this minght be associated with the preferred orientation of triplet spins in the space of $\textbf{H}_{a,b,c}$, and thus may be indicative of the order parameter symmetry. Complementary thermodynamic measurements – although challenging under such extreme magnetic fields – would be instructive for shedding further light on the remarkable quantum critical phenomena at play in UTe2.
@@
Figure 2: Suppression of the first-order metamagnetic transition towards zero temperature. (a) Contactless resistivity versus field, measured by the PDO technique (see Appendix A). The sample was mounted at a fixed angle of 17°$\degree$ from $b$ towards $c$, and subsequently rotated towards $a$. At a tilt of 6°$\degree$ away from the $bc$ plane the metamagnetic (MM) transition is still sharp and first-order-like; however, by $\theta_{a}$ = 18°$\degree$ it has markedly softened, and at 20°$\degree$ the quantum critical line (QCL) has been passed. (b) Simulation of magnetization for rotations in the same plane as panel (a), calculated from Eqn. 2 showing how the sharp jump in magnetization is suppressed upon passing the QCL. (c) PDO data for the $b-a$ ([010]–[100], top) and $b-ac$ ([010]–[101], bottom) rotation planes. All curves to $<$ 60 T were measured in a single coil magnet in Wuhan at $T=$ 4.2 K, while those to $>$ 60 T were measured in a dual-coil magnet in Dresden at 1.7 K; note that the Dresden data have been rescaled for ease of comparison, while the nature of the dual-coil magnet system introduces a small artifact at 39 T (see Fig. 15 for details). Like in panel (a), the MM transition is suppressed towards zero temperature as the QCL is approached. (d) Temperature evolution of the MM transition for H $\parallel b$ (top) and at $\theta_{b-ac}=$ 42°$\degree$ (bottom); derivatives are plotted in panel (e). The 9.0 K transition for H $\parallel b$ is considerably sharper than the 1.5 K curve at $\theta_{b-ac}$ = 42°$\degree$, which lies close to the QCL.
Figure 3: Mapping the quantum critical line in three dimensions. (a) PDO data measured in Dresden rotating from $\theta_{b-c}$ = 7°$\degree$ towards $a$ at 1.7 K. Within a one degree inclination from $\theta_{a}$ = 17°$\degree$ in this plane, the QCL is abruptly crossed. (b) PDO data of the MM transition in the $b-c$ plane, measured up to 57 T in Wuhan and up to 80 T in Dresden. The Dresden data has been rescaled by a factor of 1.1 to be comparable with that from Wuhan (see Fig. 18). (c) Positive quadrant of $H_{a}-H_{c}$ showing how the QCEPs were identified from our rotational measurements in high magnetic fields. The size of orange points corresponds to the size of the MM transition.
@@
Figure 4: Spillover of SC3 outside the polarised paramagnetic phase. (a) Electrical resistivity $\rho(H)$ at successive magnetic field tilt angles in the $b-ac$ rotation plane at $T=$ 0.4 K. At $\theta_{b-ac}=$ 17.5°$\degree$ zero resistance is observed over the field interval of 0 T $\leq\mu_{0}H\leq$ 41 T as each of SC1, SC2 and SC3 are successively accessed. (b) The phase landscape of UTe2 plotted radially in $\theta_{b-ac}$. An extended region of the onset of SC3 is observed outside the PPM state, identified by the turning over from positively to negatively sloped $\rho(H)$ marked by the black dashed line in panel (a). The red circles defining the PPM boundary were measured above $T_{c}^{SC3}$. (c) The evolution of the SC1, SC2, SC3 and PPM phases for $\theta_{b-ac}=$ 0°$\degree$ (left) and $\theta_{b-ac}=$ 35°$\degree$ (right). At $\theta_{b-ac}=$ 35°$\degree$ the melting of PPM is not observed at 5 K (marked by a cross – see Fig. 6). This shows that SC3 extends outside the PPM state for H tilted close to the QCL. (d) A comparison of the temperature sensitivity of SC3 at $\theta_{b-ac}=$ 18°$\degree$ (left) and $\theta_{b-ac}=$ 35°$\degree$ (right). On the edge of the SC3 dome at $\theta_{b-ac}=$ 18°$\degree$ the $T_{\text{c}}$ is only $\approx$ 0.5 K, whereas at $\theta_{b-ac}=$ 35°$\degree$ SC3 is much more robust to elevated temperatures. (e) Effective resistivity at $\theta_{b-ac}=$ 35°$\degree$ for incremental excitation currents as indicated. (f) Log-log plot of voltage $V$ versus current $I$ from 0.4-1.8 K at $\theta_{b-ac}=$ 35°$\degree$. Ohmic $I-V$ characteristics are observed in the normal state at $\mu_{0}H=$ 25 T for all currents, whereas at $\mu_{0}H=$ 41 T the $I-V$ behavior is highly non-Ohmic. We therefore identify the turning over of the $\rho(H)$ traces in (a) as marking the onset of SC3, characterized by this strongly non-linear relationship between $I$ and $V$.
@@
Figure 5: Anomalous $I-V$ characteristics of SC3. The exponent $\alpha$ extracted from fitting $V\propto I^{\alpha}$ to isothermal field sweeps at incremental currents, plotted as a function of field at incremental temperatures (left) and as a function of temperature for incremental fields (right). The raw data from which this figure is constructed are plotted in Figure 11.
@@
Figure 6: (a) $\rho(H)$ for H $\parallel b$ at incremental temperatures, with corresponding $\nicefrac{{\partial\rho}}{{\partial H}}$ in (b). The inset shows the derivative of the 28 K curve. (c) $\rho(H)$ at $\theta_{b-ac}$ = 35°$\degree$, with corresponding $\nicefrac{{\partial\rho}}{{\partial H}}$ in (d). These data were used to construct Fig. 4c. In all panels, arrows mark zero $\nicefrac{{\partial\rho}}{{\partial H}}$, identifying the location of the metamagnetic transition. Importantly, zero resistance is observed at $\mu_{0}H\approx$ 40 T at low temperatures for $\theta_{b-ac}=$ 35°$\degree$ even though $\nicefrac{{\partial\rho}}{{\partial H}}$ rises monotonically at $T$ = 5 K, indicating that in the low temperature limit for this orientation the metamagnetic transition field is located significantly above the maximum accessed field strength of 41.5 T for these measurements – even though zero resistance is observed at low temperatures. This shows that the SC3 phase has a region of spillover outside the PPM state.

### v3

@@ II.3 Landau theory of metamagnetic transitions in three dimensions
The SC3 phase of UTe2, which persists to $\mu_{0}H\gtrapprox$ 70 T despite its relatively modest $T_{c}$ of 2.4 K [50], has posed a formidable theoretical challenge to our understanding of high-$H$ superconductivity [17, 18]. As prior high-$H$ measurements in the $b-c$ plane only observed SC3 coexisting alongside the PPM state, the field-induced compensation of magnetic exchange [51] was proposed as a possible microscopic origin [33]. However, our finding of the zero resistance SC3 state extending outside the PPM (Fig. 5) – accompanied by an extended field region of the onset of SC3 even further away from the MM phase transition line upon approaching the QCL – argues strongly against this scenario.
@@
where $i,j,k=a,b,c$ and $\chi^{-1}_{i},\beta_{ij}$, and $\delta_{ijk}$ are model parameters (see Appendix B[52, 53, 54]. The MM phase transition occurs when the free energy has two local minima at $\mathbf{M}=\textbf{M}_{0}$ and $\mathbf{M}=\mathbf{M}_{*}$ corresponding to the normal paramagnetic and PPM phases, respectively. With $\beta_{bb}<0$, the $H_{b}$ component of the external field tends to lower the minimum at $\mathbf{M}_{*}$ faster than the minimum at $\mathbf{M}_{0}$, resulting in the MM transition when $\mathcal{F}(\mathbf{M}_{*})=\mathcal{F}(\mathbf{M}_{0})$, which generically defines a 2D surface in the 3D space of $\mathbf{H}$.
@@
In general, the model parameters are temperature dependent, such that as temperature increases the two minima merge at a CEP. For simplicity, here we focus on the low temperature limit and consider the effects of $H_{a}$ and $H_{c}$. Due to the mixed terms $\beta_{ij}$ and $\delta_{ijk}$, both $H_{a}$ and $H_{c}$ (which induce $M_{a}$ and $M_{c}$) in effect indirectly modify the $\chi_{b}$ and $\beta_{bb}$ coefficients and so act in a similar way to the temperature. As a result, large $H_{a}$ and $H_{c}$ tend to move $\mathbf{M}_{0}$ and $\mathbf{M}_{*}$ towards each other, until they merge at the critical end line (at finite temperature) or the QCL (at zero temperature) bounding the MM phase transition surface.
@@
We confirm this general picture by numerically minimizing the free energy of Eqn. (2), determining the MM transition surface and the QCL as a function of H, as shown in Figure 9. With the given parameters (see Appendix B) the model matches the experimental data from Section II.1 reasonably well and captures the main qualitative features. Assuming SC3 is driven by the fluctuations in the vicinity of the QCL, the shape of the QCL is qualitatively close to the observed field direction dependence of SC3, with orbital pair-breaking and spin effects potentially also relevant. The QCL fluctuation scenario is further supported by the fact that, analogous to the case of ferromagnetic uranium-based superconductors such as URhGe [12, 13, 53], SC3 should be expected to occur on both sides of the QCL as we observe experimentally (Fig. 5).
III Discussion
In combination, our high-$H$ measurements of UTe2 presented in Section II (i) identify that an extended line of QCEPs bounds the surface of first-order MM transitions in three-dimensional H-space at the remarkably high magnetic field scale of over 70 T, (ii) anchor high-$H$-induced superconductivity in UTe2 close to this boundary, and (iii) reveal that in a narrow angular range close to the QCL the onset of the SC3 state extends down to fields that lie far below the MM transition field.
We note that the concept of a QCL has previously been considered in holographic theoretical studies of quantum criticality [55, 56]. Additionally, extended lines of quantum criticality have been reported in experimental investigations of the pressure and doping phase diagrams of several cerium-based heavy fermion superconductors [57, 58, 59, 60]. For example, in the case of CeCoIn5, the QCL of this material is a function of pressure $p$ and $H$ [59]. The QCL connects a series of quantum critical points located at the edge of antiferromagnetic order, which trace out a smooth curve when plotted in two-dimensional $p-H$ space. By contrast, in the present study we find that the sensitivity of the termination of UTe2’s PPM state to each of the three orthogonal directions of H leads to a three-dimensional QCL at ambient pressure. Given that the field scale of both SC3 and the MM phase boundary have been reported to decrease with increasing $p$ [54, 61], it is possible that under compression the QCL of UTe2 would be extruded to form a quantum critical surface in the four-dimensional space of $p$-H. Further high-$H$ measurements under pressure are required to test this hypothesis.
Our findings here – of $H$-induced superconductivity in UTe2 being located in proximity to QCEPs of a first-order MM phase boundary – mirror prior studies on the related compound URhGe [11, 12, 62, 13, 63]. This material possesses a ferromagnetic groundstate, with moments oriented along the $c$-axis. For H applied along the $b$ direction, a metamagnetic transition occurs at $\mu_{0}H_{b}\approx$ 12 T, which can be interpreted as a rotational transition of the orientation of magnetic moments from $\parallel c$ to $\parallel b$ [11], with a corresponding change of the easy magnetic axis [63]. $H$-induced superconductivity is observed at low temperatures over the interval of 8 T $\lessapprox\mu_{0}H_{b}\lessapprox$ 13 T [11]. This magnetic field domain lies predominantly within the ferromagnetic phase, with a small portion of the superconductivity spilling out to the other side of the MM transition. Applying an $a$-axis magnetic field component has very little effect on the $H$-induced superconductivity, which can survive up to (at least) $\mu_{0}\absolutevalue{\textbf{H}}=$ 32 T [62] (where here |$\absolutevalue{\textbf{H}}\equiv\sqrt{H_{a}^{2}+H_{b}^{2}}$). However, at $\mu_{0}H_{b}=$ 12 T, just a small $c$-axis field component of $\mu_{0}H_{c}<$ 2 T quickly destroys the $H$-induced superconductivity, as the QCEP of the first-order MM phase boundary is crossed [12, 62, 63]. Nuclear magnetic resonance (NMR) measurements have provided microscopic evidence that the $H$-induced superconductivity is mediated by quantum critical magnetic fluctuations stemming from the suppression of the first-order MM phase boundary to zero temperature [64, 65].
There are numerous similarities between URhGe and our findings here on UTe2. The MM transition in UTe2 also marks a change in the easy axis – in this case, from $a$ to $b$ [35]. Both materials possess $H$-induced superconductivity in proximity to the QCEPs of a first-order MM phase boundary, the CEP of which may be tuned to absolute zero by the tilt angle of the magnetic field. While the $H$-induced superconductivity of URhGe lies predominantly within the ferromagnetic state, extending a little to higher $H$ on the other side of the MM transition, in UTe2 SC3 resides predominantly within the PPM state, extending a little to lower $H$ on the other side of the MM transition (Fig. 5). The key differences between the two materials are the magnetic ground state properties, the associated magnetic and superconducting energy scales, and the dimensionality of their quantum critical phase boundaries.
At first glance, perhaps the most striking contrast between these two materials is the remarkably high magnetic field scales of the metamagnetic and associated quantum critical phenomena underpinning high-$H$-induced superconductivity in UTe2. A commonality shared by both compounds is that their $H$-induced superconducting states possess higher maximal $T_{c}$ values than is observed for the superconductive phases at ambient magnetic field [11]. However, at 2.4 K [50], the maximal $T_{c}$ of UTe2 is markedly higher than the comparable value of 0.4 K in the case of URhGe. The location of $\mu_{0}H_{m}$ at 34 T is also considerably higher than URhGe’s value of 12 T. Regarding the magnetic properties, at low $H$ UTe2 lacks long-range magnetic order; however, it exhibits clear signatures of antiferromagnetic (high wavevector $Q$) correlations [66, 67, 68]. It is not clear if these fluctuations persist up to the high field strengths at which the QCL and SC3 are found. Given that the quantum critical phase boundary of UTe2 is sensitive to all three orthogonal spatial magnetic field components – unlike URhGe, in which only two dimensions are relevant – it is not inconceivable that the SC3 state might be mediated by a combination of low and high $Q$ fluctuations. For H $\parallel b$, NMR measurements of UTe2 have observed strong ferromagnetic-like longitudinal spin fluctuations for $\mu_{0}H_{b}\gtrapprox$ 15 T, which were found to diverge as $H\rightarrow H_{m}$ [69]. These findings appear to concur with an analysis of the magnetic entropy landscape upon approaching the MM transition, in which strong $Q=0$ fluctuations were discerned [70].
Our discovery of the presence of an extended line of quantum criticality at high magnetic fields in UTe2 provides a natural explanation for how this material is able to manifest remarkably high-$H$-induced superconductivity. However, understanding the exact shape of the SC3 pocket in H-space remains an outstanding puzzle. In Fig. 3b we schematically depicted the SC3 domain with purple shading, guided by measurements from this study combined with those reported in refs. [33, 36]. (See Figs. 3b & 8 for the data that constrain this shading, and Fig. 21 for a comparison of our dataset with that of ref. [36].) While for $H_{a}\neq$ 0 the QCL closely bounds the highest-$H$ locations of SC3 observed by our measurements and those of ref. [36], for $H_{a}=0$ (i.e. in the $b-c$ plane) there is a large discrepancy between the termination of the SC3 state and the QCL. In Fig. 20 we show how the diminution of the MM transition in the $b-c$ plane extrapolates to indicate the QCEP is located at $\mu_{0}H\approx$ 77 T in this plane, which would correspond to an angle $\theta_{b-c}>$ 60°$\degree$ (Fig. 18). However, multiple studies by independent groups indicate that the maximal upper critical field and $T_{c}$ of SC3 in the $b-c$ plane lies in the range 30°$\degree<\theta_{b-c}<40\degree$, and appears to terminate at $\theta_{b-c}\approx 45\degree$ (see Fig. 19[33, 49, 50]. This is despite the MM transition still giving a sharp discontinuity in $\Delta f_{\text{PDO}}$ at $\theta_{b-c}=60\degree$ (Fig. 20).
This hints at further subtleties in the exotic high-$H$ physical properties of UTe2. One possible explanation concerns the d-vector of SC3 (if it is odd-parity in character, which appears highly likely). When Cooper pairs possess non-zero spin, the gap structure of the superconducting order parameter can set strong constraints on their preferred orientation with respect to H, which may induce complex magnetic field strength-angle-temperature phase diagrams [10, 71]. It has recently been proposed that the toroidal shape of the SC3 pocket might be explained by a multi-component non-unitary odd-parity order parameter, with the Cooper pairs possessing a finite orbital momentum predominantly oriented along the $c$-axis [36]. Given that $c$ is the hard magnetic axis of UTe2 at high $H$ [35], such a scenario appears plausible. Alternatively, perhaps there is no QCEP for $H_{a}=0$ (i.e. there should be no open green stars in Fig. 3), but instead the QCL asymptotically bends towards the $H_{a}=0$ axis as $H_{c}\rightarrow\infty$ (i.e. as $\theta_{b-c}\rightarrow 90\degree$). If true, such a scenario would again be analogous to the case of URhGe, in which $a$ is the hard axis and thus $H_{a}$ has a negligible effect on the $H$-induced superconductivity, which would presumably persist up to arbitrarily large values of $H_{a}$ until the orbital depairing threshold is reached. Further measurements of UTe2’s SC3 state in proximity to the QCL (at $H_{a}\neq 0$) by bulk-sensitive thermodynamic probes to track departures from non-Fermi liquid phenomenology, and by NMR of the spin relaxation rates upon approaching the QCL – although challenging due to the high magnetic field scales and practical difficulties in sample orientation – would be valuable for shedding further light onto the microscopic properties of the remarkable three-dimensional quantum critical phase boundary manifested by this material.
In summary, here we have presented a detailed mapping of the high magnetic field phase landscape of UTe2. We tracked the profile of this material’s $H$-induced first-order metamagnetic transition surface as functions of temperature and of magnetic field tilt angle. We find that for rotations of H away from the crystallographic $b$-axis, the critical end point of the metamagnetic transition is smoothly suppressed towards zero temperature. Because the metamagnetic transition surface extends throughout the three dimensions of magnetic field space, this implies that it is bounded by a continuous ring of quantum critical end points, the locus of which forms an extended line of quantum criticality. For H oriented close to this quantum critical line, the high-$H$ SC3 superconductive state is observed on both sides of the metamagnetic transition, similar to the case of $H$-induced superconductivity in the related material URhGe [11, 12]. Magnetic fluctuations emanating from this three-dimensional quantum critical phase boundary therefore appear the likely source of electronic pairing mediating the formation of $H$-induced superconductivity in UTe2.
Figure 1: Suppression of a first-order metamagnetic transition towards zero temperature. (a) Schematic temperature $T$ versus magnetic field $H$ depiction of the critical endpoint (CEP) of a first-order metamagnetic transition. The solid orange line marks the first-order phase boundary, while the dashed (Widom) line indicates crossover behavior at higher temperatures. (b) The modulation of a non-thermal tuning parameter – such as a magnetic field – can suppress a CEP to lower temperatures. When it reaches absolute zero, it becomes a quantum critical endpoint (QCEP). (c) Isothermal contactless resistivity measurements of UTe2, obtained by the PDO technique (see Appendix A), offset for clarity. The sample was mounted at a fixed angle of 17°$\degree$ from $b$ towards $c$, and subsequently rotated towards $a$ (see Fig. 11 for a description of this rotation procedure). At a tilt of 6°$\degree$ away from the $bc$ plane the metamagnetic (MM) transition is sharp and first-order-like. However, at $\theta_{a}$ = 18°$\degree$ it has markedly softened and by 20°$\degree$ the QCEP has been passed. Dashed lines and green star give a visual depiction of how the tilt angle of the magnetic field tunes the CEP to a QCEP. (d) Simulation of magnetization for rotations in the same plane as panel (c), calculated from Eqn. 2 showing how the sharp jump in magnetization is suppressed upon approaching a QCEP. (e) Temperature evolution of the MM transition for H $\parallel b$ (top) and at $\theta_{b-ac}=$ 42°$\degree$ (bottom); offset derivatives are plotted in panel (f). The 9.0 K transition for H $\parallel b$ is considerably sharper than the 1.5 K curve at $\theta_{b-ac}$ = 42°$\degree$, indicating that the CEP is located at a much lower temperature at this angle
Figure 2: Mapping quantum criticality in three dimensions. PDO data for the $b-a$ ([010]–[100]), $b-ac$ ([010]–[101]) and $b-c$ ([010]–[001]) rotation planes, along with rotations towards $a$ at a fixed inclination of 7°$\degree$ from $b$ to $c$. All curves to $<$ 60 T were measured in a single coil magnet in Wuhan at $T=$ 4.2 K, while those to $>$ 60 T were measured in a dual-coil magnet in Dresden at 1.7 K; note that the Dresden data have been rescaled for ease of comparison, while the nature of the dual-coil magnet system introduces a small artifact at 39 T (see Fig. 17 for details). For each of these rotation planes, the size of the MM transition decreases as H rotates further from $b$, as expected for the behavior depicted in Fig. 1b. This is especially pronounced in the bottom panel, where a small change in inclination from 17°$\degree$ to 18°$\degree$ abruptly crosses the CEP at this relatively low measurement temperature of 1.7 K.
@@ III Discussion
Figure 4: Determination of the quantum critical line (QCL). Here we plot the positive quadrant of $H_{a}-H_{c}$, viewed looking along the $H_{b}$ direction, to show how the QCEPs were identified from our rotational magnetoconductance measurements in high magnetic fields. The size of orange points corresponds to the size of the MM transition at that location. The green stars indicate the extrapolation of diminishing MM transition size, to yield the position of the QCEP in each rotation plane. The locus of QCEPs thereby forms the QCL.
@@
Figure 5: Spillover of SC3 outside the polarized paramagnetic phase. (a) Electrical resistivity $\rho(H)$ at successive magnetic field tilt angles in the $b-ac$ rotation plane at $T=$ 0.4 K. At $\theta_{b-ac}=$ 17.5°$\degree$ zero resistance is observed over the field interval of 0 T $\leq\mu_{0}H\leq$ 41 T as each of SC1, SC2 and SC3 are successively accessed. For ease of presentation only selected curves are plotted – see Fig. 12 for the full angular dataset. (b) The phase landscape of UTe2 plotted radially in $\theta_{b-ac}$. An extended region of the onset of SC3 is observed outside the PPM state, identified by the turning over from positively to negatively sloped $\rho(H)$ marked by the black dashed line in panel (a). The red circles defining the PPM boundary were measured above $T_{c}^{SC3}$ (see Fig. 12). (c) The evolution of the SC1, SC2, SC3 and PPM phases for $\theta_{b-ac}=$ 0°$\degree$ (left) and $\theta_{b-ac}=$ 35°$\degree$ (right). At $\theta_{b-ac}=$ 35°$\degree$ the MM transition is not observed at 5 K (marked by a cross – see Fig. 7). This shows that SC3 extends outside the PPM state for H tilted close to the QCL. (d) A comparison of the temperature sensitivity of SC3 at $\theta_{b-ac}=$ 18°$\degree$ (left) and $\theta_{b-ac}=$ 35°$\degree$ (right). On the edge of the SC3 dome at $\theta_{b-ac}=$ 18°$\degree$ the $T_{\text{c}}$ is only $\approx$ 0.5 K, whereas at $\theta_{b-ac}=$ 35°$\degree$ SC3 is much more robust to elevated temperatures. (e) Effective resistivity at $\theta_{b-ac}=$ 35°$\degree$ for incremental excitation current densities as indicated. Data at other temperatures are presented in Fig. 13. (f) Log-log plot of voltage $V$ versus current density $J$ from 0.4-1.8 K at $\theta_{b-ac}=$ 35°$\degree$. Ohmic $J-V$ characteristics are observed in the normal state at $\mu_{0}H=$ 25 T for all currents, whereas at $\mu_{0}H=$ 41 T the $J-V$ behavior is highly non-Ohmic. We therefore identify the turning over of the $\rho(H)$ traces in (a) as marking the onset of SC3, characterized by this strongly non-linear relationship between $J$ and $V$.
@@
Figure 6: Anomalous current–voltage characteristics of SC3. The exponent $\alpha$ extracted from fitting $V\propto J^{\alpha}$ to isothermal field sweeps at incremental currents, plotted as a function of field at incremental temperatures (left) and as a function of temperature for incremental fields (right). The raw data from which this figure is constructed are plotted in Figure 13. This analysis indicates that non-Ohmic electrical transport properties – indicative of the onset of the SC3 state – onset from $\mu_{0}H\approx 30$ T for $\theta_{b-ac}=$ 35°$\degree$, coinciding with the maximum in $\rho(H)$ (in the low $J$ limit) for this orientation.
@@
Figure 7: (a) $\rho(H)$ for H $\parallel b$ at incremental temperatures, with corresponding $\nicefrac{{\partial\rho}}{{\partial H}}$ in (b). The inset shows the derivative of the 28 K curve. (c) $\rho(H)$ at $\theta_{b-ac}$ = 35°$\degree$, with corresponding $\nicefrac{{\partial\rho}}{{\partial H}}$ in (d). These data were used to construct Fig. 5c. In all panels, arrows mark zero $\nicefrac{{\partial\rho}}{{\partial H}}$, identifying the location of the metamagnetic transition. Importantly, zero resistance is observed at $\mu_{0}H\approx$ 40 T at low temperatures for $\theta_{b-ac}=$ 35°$\degree$ even though $\nicefrac{{\partial\rho}}{{\partial H}}$ rises monotonically at $T$ = 5 K, indicating that in the low temperature limit for this orientation the metamagnetic transition field is located significantly above the maximum accessed field strength of 41.5 T for these measurements – even though zero resistance is observed at low temperatures. This shows that, in proximity to the QCL, the SC3 phase has a region of spillover outside the PPM state.